92g:20006 20C05 16S34 20C20

Siegel, Stephen ^{(}1-CHI^{)}

Projective modules for A_{9} in characteristic three. Comm. Algebra 19 (1991), no. 11, 3099–3117. The author determines the Loewy structures of the indecomposable projective modules in characteristic 3 for the alternating group of degree 9. There are three blocks: one of defect zero, one of defect 1, and the principal block. All indecomposable projective modules in the principal block have Loewy length 7.

Burkhard Kulshammer (D-AGSB)

\begin{matrix}c \\ ◯ \end{matrix} Copyright American Mathematical Society 1992, 1996